A card is drawn from a deck of 20 cards (all the cards numbered 1 to 5 in a standard check) and the random variable W gives the number on the card. Construct the probability mass function and its corresponding histogram.


1
Expert's answer
2022-02-27T15:12:06-0500

"P(W=1)=P(W=2)=P(W=3)=P(W=4)=P(W=5)=\frac{4}{20}=\frac{1}{4}."